### v1

# ReMoD: Rethinking Modality Contribution in Multimodal Stance Detection via Dual Reasoning

Abstract
Multimodal Stance Detection (MSD) is a crucial task for understanding public opinion on social media. Existing work simply fuses information from various modalities to learn stance representations, overlooking the varying contributions of stance expression from different modalities. Therefore, stance misunderstanding noises may be drawn into the stance learning process due to the risk of learning errors by rough modality combination. To address this, we get inspiration from the dual-process theory of human cognition and propose ReMoD, a framework that Rethinks Modality contribution of stance expression through a Dual-reasoning paradigm. ReMoD integrates experience-driven intuitive reasoning to capture initial stance cues with deliberate reflective reasoning to adjust for modality biases, refine stance judgments, and thereby dynamically weight modality contributions based on their actual expressive power for the target stance. Specifically, the intuitive stage queries the Modality Experience Pool (MEP) and Semantic Experience Pool (SEP) to form an initial stance hypothesis, prioritizing historically impactful modalities. This hypothesis is then refined in the reflective stage via two reasoning chains: Modality-CoT updates MEP with adaptive fusion strategies to amplify relevant modalities, while Semantic-CoT refines SEP with deeper contextual insights of stance semantics. These dual experience structures are continuously refined during training and recalled at inference to guide robust and context-aware stance decisions. Extensive experiments on the public MMSD benchmark demonstrate that our ReMoD significantly outperforms most baseline models and exhibits strong generalization capabilities.
Introduction
The proliferation of social platforms supporting multimodal content, such as text or images (8-dar2025social; 9-bansal2024hybrid), has made Multimodal Stance Detection (MSD) a critical task for understanding public opinion (1-liang2024multi; 2-niu2024multimodal). MSD aims to identify the expressed attitude toward a specific target by integrating both textual and visual information, enabling a more comprehensive analysis than purely textual approaches (5-augenstein2016stance; 6-liang2022zero). This task applies widely to political discourse analysis, misinformation detection, and public health communication, making it a crucial tool for decoding complex online interactions (2-niu2024multimodal; 3-pangtey2025large).
Figure 1: Examples of different samples with various modality contributions.
Existing MSD methods focus on enhancing the capability of extracting multimodal representations (1-liang2024multi; 31-zhang2025logic). However, these methods simply fuse information from different modalities to learn stance representations, overlooking the fact that the contributions of different modalities in some examples to the final stance decision can vary. Figure 1 demonstrates how modality contributions vary, illustrating two issues not considered in existing MSD tasks. First, the informative value of individual modalities is highly context-dependent, as seen in Figure 1(a) and (b), where a modality may be ambiguous or irrelevant in isolation, providing little evidential support for stance detection. Models that fail to determine modal contributions are thus susceptible to being skewed by such low-quality inputs. Second, the semantic relationship between modalities is not fixed but dynamic, ranging from complementary to contradictory as shown in Figure 1(c) and (d). Simple feature alignment is ill-equipped to navigate these nuanced interactions. Therefore, when encountering modalities of varying contributions, these methods tend to inappropriately fuse low-quality modal information with high-confidence judgments, thereby contaminating the entire detection process.
To bridge this gap, we introduce ReMoD, a novel framework for multimodal stance detection that Rethinks Modality contribution of stance expression by integrating a Dual-reasoning paradigm inspired by the dual-process theory of human cognition. ReMoD integrates two core reasoning stages to dynamically weight modality contributions based on their actual expressive power for the target stance. First, an experience-driven intuitive reasoning stage instantly generates an initial stance hypothesis by querying the Modality Experience Pool (MEP) and Semantic Experience Pool (SEP), prioritizing historically impactful modality. This initial hypothesis is then passed to a deliberate reflective reasoning stage for critical scrutiny via two complementary reasoning chains. The Modality Chain of Thought (Modality-CoT) reflects on the initial prediction to formulate adaptive fusion strategies, dynamically learning how to weigh different modalities. These strategies are then used to update the MEP, creating a self-improving loop for modality handling. Concurrently, the Semantic Chain of Thought (Semantic-CoT) analyzes the content’s contextual nuances to distill deeper contextual insights, which are updated to SEP. During training, ReMoD continuously refines these dual experience pools, and at inference, it recalls this accumulated experience to make robust, context-aware decisions. Our main contributions are as follows:
  • We propose ReMoD, a novel framework that pioneers a dual-reasoning paradigm for multimodal stance detection. By synergizing experience-driven intuitive reasoning with deliberate reflective reasoning, it realizes a human-like adaptive reasoning process tailored to complex multimodal scenarios, addressing the oversight of dynamic modality contributions in existing methods.
  • We design the Modality Experience Pool and Semantic Experience Pool integrated with a modality-synergistic experience retrieval, which dynamically accumulates modality contribution patterns and semantic insights.
  • Extensive experiments conducted on the public MMSD benchmark demonstrate the superiority of our approach. ReMoD significantly outperforms most baseline models and exhibits strong generalization capabilities.
Figure 2: Overflow of our ReMoD framework for multimodal stance detection.
Related Works
LLM-based Stance Detection
The design of modern stance detection frameworks has been significantly influenced by LLMs. Early approaches focused on harnessing their zero-shot and few-shot capabilities, using direct prompting and Chain-of-Thought (CoT) to elicit stance predictions without labeled data or fine-tuning (13-zhang2023investigating; 14-zhang2022would). Other works have pursued fine-tuning, for instance, by fusing retrieved background knowledge with input pairs to improve learning through backpropagation (11-li2023stance). Subsequent research has sought to further unleash LLM reasoning. Innovations include integrating multi-perspective insights via CoT and prompt-tuning (15-ding2024cross), formalizing reasoning with first-order logic rules for enhanced interpretability (16-zhang2024knowledge), and architecting multi-agent systems where LLMs collaborate for detection (12-lan2024stance; 19-wang2024deem). However, these methods are primarily developed for pure textual stance detection and have not yet explored the integration of multimodal information such as images.
Multimodal Stance Detection
The proliferation of social media and video platforms has driven the rise of multimodal content (text, image) in forms like memes and reels (8-dar2025social; 9-bansal2024hybrid), which convey implicit opinions through diverse signals, posing challenges for stance detection. Modeling semantically heterogeneous cross-modal interactions remains difficult for traditional models, whereas transformer architectures and LLMs excel here via self-attention and transfer learning. Recent approaches increasingly adopt multimodal fusion strategies to improve stance classification. For example, 2-niu2024multimodal proposed MmMtCSD and MLLM-SD for text-image pairs in multi-turn dialogues; 23-kuo2024advancing fused textual, visual, and social signals in political discourse using S-BERT, BEiT, and GraphSAGE. SERN (24-xie2021sern) enhanced fake news detection via sentence-guided visual attention; CT-TN (25-khiabani2023few) and 26-barel2025acquired addressed few-shot/noisy settings with graph-based cues; 27-yao2023end extended MSD to fact-checking via web evidence integration. However, these approaches typically employ rigid fusion strategies, overlooking the dynamic interplay and context-dependent reliability of modalities, which makes them vulnerable to noisy or contradictory information.
Method
This section introduces our novel ReMoD framework for MSD in detail. Drawing inspiration from the dual-process theory of human cognition (47-kahneman2011fast), a psychological model that posits two distinct modes of thought. One is an intuitive, fast ‘System 1’ that relies on heuristics and prior experiences to make rapid judgments. The other is a deliberate, slow ‘System 2’ that engages in analytical reasoning to refine and validate these initial assessments. Our ReMoD framework operationalizes this cognitive paradigm for multimodal stance detection.
Given a text-image pair $(u,v)$, the aim of ReMoD is to identify the stance label $y$, towards a specific target $t$. The architecture of our ReMoD is depicted in Figure 2, which consists of three primary components: 1) Multimodal Knowledge Perception, which extracts comprehensive multimodal information and captures relevant external knowledge. 2) Experience-Driven Intuitive Reasoning (mirroring System 1), which queries the Modality Experience Pool to instantly form an initial stance hypothesis; and 3) Deliberate Reflective Reasoning (emulating System 2), which employs complementary reasoning chains to scrutinize the initial hypothesis and continuously refine the dual experience pools.
Multimodal Knowledge Perception
This initial stage constructs a rich semantic foundation by extracting and refining knowledge from both textual and visual modalities. The goal is to capture not only high-level conceptual information but also grounded, fine-grained evidence, which is crucial for nuanced reasoning.
Textual Knowledge Attainment.
For the text input $u$, we perform a three-step knowledge extraction process. First, we employ a Multimodal Large Language Model (MLLM) with strategically designed prompts to identify a set of key semantic entities, denoted as $\mathcal{E}_{u}$. To ground these entities in external world knowledge, we retrieve corresponding background information for each entity $e\in\mathcal{E}_{u}$ from the Wikipedia knowledge base. As the raw retrieved knowledge is often verbose and contains noisy information, we leverage the MLLM again to distill it into a concise, semantically-rich summary. This final summary, denoted as the text knowledge $\mathcal{K}_{u}$, serves as the knowledge for the text modality.
Visual Knowledge and Feature Grounding.
For the image input $v$, we capture both high-level semantic concepts and low-level visual information. First, to extract semantic concepts, an MLLM generates an image caption, identifies a set of visual entities $\mathcal{E}_{v}$ from the caption. These entities are then used to query Wikipedia, and the resulting information is summarized into a condensed visual knowledge $\mathcal{K}_{v}$. However, knowledge alone is insufficient to capture the visual manifestation of these concepts. To address this, we ground the semantic entities in the image’s pixel space. We leverage the identified entity set $\mathcal{E}_{v}$ to guide the open-vocabulary segmentation model SAM (29-ravi2024sam) in precisely segmenting the image regions corresponding to each entity, yielding a set of entity-aligned visual information $\mathcal{V}$ where each element is a segmented image directly showcasing the visual appearance of a specific semantic entity.
$\mathcal{V}=\{\bar{v}_{i}\}^{N}_{i=1}=\text{SegModel}(v,\mathcal{E}_{v})$ (1)
where $N$ indicates the number of sub-images.
Experience-Driven Intuitive Reasoning
This stage operationalizes the ‘fast thinking’ of our dual-reasoning paradigm. It simulates rapid, intuitive judgment by leveraging experience pools accumulated during training to quickly formulate an initial stance hypothesis. The process consists of two main steps: modality-synergistic experience retrieval and experience-driven stance prediction.
Modality-Synergistic Experience Retrieval.
To retrieve relevant past experiences, we introduce a modality-synergistic retrieval mechanism. Unlike conventional text-only systems, our approach leverages both text and visual information to query the modality experience pool and semantic experience pool. Each pool stores a collection of experiences $\{E_{j}\}$, where each experience is a key-value pair. The key is a pre-computed bimodal embedding $(k_{u}^{j},k_{\mathcal{V}}^{j})$, representing the experience’s core features, while the value contains the associated experiential data.
The retrieval process commences by encoding the current text input text $u$ and sub-images $\mathcal{V}$ into their respective query vectors, $q_{u}$ and $q_{\mathcal{V}}$:
$q_{u}=\text{BGE-M3}(u),~q_{\mathcal{V}}=\frac{1}{N}\sum^{N}_{i=1}\text{ImageEncoder}(\bar{v}_{i})$ (2)
where $q_{u}$ is generated by the BGE-M3 text encoder (46-chen2024bge), and $q_{\mathcal{V}}$ is produced by the image encoder of the MLLM. The stored key vectors $k_{u}^{j}$ and $k^{j}_{\mathcal{V}}$ for each Experience $E_{j}$ are computed analogously as $q_{u}$ and $q_{\mathcal{V}}$.
With the query vectors established, we measure their similarity to the key vectors of each stored experience $E_{j}$ by computing unimodal relevance scores for text $S_{u}$ and vision $S_{v}$ using cosine similarity:
$S_{u}(q_{u},k_{u}^{j})=\frac{q_{u}\cdot k_{u}^{j}}{\|q_{u}\|\|k_{u}^{j}\|},~S_{\mathcal{V}}(q_{\mathcal{V}},k_{\mathcal{V}}^{j})=\frac{q_{\mathcal{V}}\cdot k_{\mathcal{V}}^{j}}{\|q_{\mathcal{V}}\|\|k_{\mathcal{V}}^{j}\|}$ (3)
To obtain a holistic measure of relevance, we fuse these scores into a unified Modality Relevance Score (MRS), denoted as $S(E_{j})$, via a weighted linear combination:
$S(E_{j})=\alpha\cdot S_{u}(q_{u},E_{j})+(1-\alpha)\cdot S_{\mathcal{V}}(q_{\mathcal{V}},E_{j})$ (4)
where $\alpha\in[0,1]$ balances the contribution of each modality. Ultimately, all experiences are ranked in descending order based on their MRS. We then filter this ranked list, selecting the top-$k$ experiences whose scores exceed a relevance threshold $\tau$. This procedure yields the final sets of retrieved experiences including the modality experience set $\{\mathcal{S}_{i}^{\text{ME}}\}^{k}_{i=1}$ and the semantic experience set $\{\mathcal{S}_{i}^{\text{SE}}\}^{k}_{i=1}$.
Experience-driven Stance Prediction.
Following the retrieval of relevant past experiences, our ReMoD agent $A_{R}$ proceeds to generate initial, intuitive stance hypotheses. We formulate a comprehensive prompt that integrates two key components: (1) text information $u$ and visual information $\mathcal{V}$; (2) the retrieved experience sets $\{\mathcal{S}^{\text{ME}}\}^{k}_{i=1}$ and $\{\mathcal{S}^{\text{SE}}\}^{k}_{i=1}$, which serve as in-context exemplars. This prompt instructs $A_{R}$ to perform a preliminary analysis from three distinct perspectives: text-only, vision-only, and combined multimodal. For each perspective, the agent generates an initial stance prediction $\hat{y}$ and a corresponding rationale $r$. This process is formally expressed as:
$(\hat{y}_{c},r_{c})=A_{R}(\text{prompt}(c,\{\mathcal{S}^{\text{ME}}\}^{k}_{i=1},\{\mathcal{S}^{\text{SE}}\}^{k}_{i=1}))$ (5)
where $c\in\{u,\mathcal{V},(u,\mathcal{V})\}$ denotes the context of evaluation: the text informationuu, the visual information $\mathcal{V}$ or the full multimodal pair $(u,\mathcal{V})$. These generated pairs $(\hat{y}_{c},r_{c})$ serve as essential intermediate reasoning traces that are scrutinized and refined during the subsequent deliberate reflective reasoning stage.
Deliberate Reflective Reasoning
Deliberate Reflective Reasoning is analogous to the ‘System 2’ slow thinking process in dual-process cognitive theory and serves as the refinement module of our framework. This process is governed by experience formulation, where novel insights are generated through structured reflection, and experience evolution, where the experience pools are refined by integrating these new insights.
Experiences Formulation.
The primary objective of experience formulation is to facilitate metacognitive reflection, enabling the agent to distill actionable experience from its performance outcomes. To this end, the ReMoD agent employs a structured, multi-step reasoning process, analogous to a CoT, to generate new insights.
First, to generate the modality experience, ReMoD evaluates the accuracy of its initial stance predictions across text-only $\hat{y}_{u}$, vision-only $\hat{y}_{\mathcal{V}}$, and combined multimodal $\hat{y}_{(u,\mathcal{V})}$ contexts, along with their corresponding rationales $r_{u}$, $r_{\mathcal{V}}$, and $r_{(u,\mathcal{V})}$. The outcomes are then distilled into a high-level modality insight $I_{m}$. This insight functions as a strategic heuristic, guiding the optimal balance of unimodal and multimodal cues to improve stance detection performance. In parallel, forming semantic experience involves ReMoD initiating deep reflection on its reasoning process, with guidance from the initial prediction outcomes of the same text-only, vision-only, multimodal contexts, and the corresponding rationales. This reflective process involves a joint analysis of the source text knowledge $\mathcal{K}_{u}$ and visual knowledge $\mathcal{K}_{v}$ to scrutinize the depth and comprehensiveness of the reasoning that led to those predictions. These reflections are subsequently distilled into a semantic insight $I_{s}$, which articulates actionable strategies for deepening semantic understanding for the task.
Experiences Evolution.
Upon generating new insights $I_{m}$ and $I_{s}$, ReMoD updates its experience pools to consolidate and refine existing experiences by first employing the modality-synergistic retrieval mechanism to retrieve the top-$k$ most relevant prior experiences from each pool. These experiences, denoted as $\{\mathcal{\bar{S}}_{i}^{\text{ME}}\}^{k}_{i=1}$ from MEP and $\{\mathcal{\bar{S}}_{i}^{\text{SE}}\}^{k}_{i=1}$ from SEP, have scores exceeding a relevance threshold $\tau$ and are selected based on the current text information $u$ and visual information $\mathcal{V}$. If no experiences meet this threshold, the new insights are encapsulated as new key-value entries and appended to their respective pools. Conversely, if relevant experiences are found, ReMoD performs targeted updates on these entries to prevent redundant memories.
This evolutionary update hinges on a fusion operation. For each retrieved modality experience, ReMoD initiates a re-reasoning process, using a prompt to synthesize the prior experience $\{\mathcal{\bar{S}}^{\text{ME}}_{i}\}^{k}_{i=1}$ with the new modality insight $I_{m}$. This generates a more comprehensive, updated modality experience pool $\mathcal{\hat{S}}^{\text{ME}}$, which then overwrites the original entry. A parallel process is executed for the semantic experiences, updating $\{\mathcal{\bar{S}}_{i}^{\text{SE}}\}^{k}_{i=1}$ with $I_{s}$ to produce new semantic experience pool $\mathcal{\hat{S}}_{i}^{\text{SE}}$. This dual-update ensures both experience pools remain synchronized and continuously refined. In essence, this mechanism enables ReMoD to build a compact and coherent knowledge base through continuous refinement, supporting efficient and scalable lifelong learning.
MODALITY METHOD MTSE MCCQ MWTWT MRUC MTWQ
DT JB CQ CA CE AC AH DF RUS UKR MOC TOC
Text-only BERT 48.25 52.04 66.57 75.62 60.85 63.05 59.24 81.53 41.25 46.80 57.77 45.91
RoBERTa 58.39 60.79 66.57 69.56 65.03 69.74 67.99 79.21 39.52 57.66 55.22 48.88
KEBERT 64.50 69.81 66.84 71.67 67.56 69.29 69.74 80.57 41.55 59.01 58.15 47.75
LLaMA2 53.23 52.67 47.40 34.89 41.95 49.09 44.32 30.21 38.84 38.54 55.31 46.51
GPT4 68.74 66.39 65.84 63.14 65.12 69.93 71.62 52.69 41.64 53.76 58.05 49.81
Vision-only ResNet 37.89 38.59 47.16 39.89 42.20 43.52 37.05 50.34 35.10 40.00 42.02 33.94
ViT 40.48 40.42 46.64 46.63 50.00 40.16 46.32 50.86 33.31 39.87 38.63 35.53
SwinT 39.89 40.43 48.80 46.30 46.99 41.02 47.39 51.32 35.01 40.89 35.03 35.47
Multimodal BERT+ViT 41.86 45.82 61.32 63.20 44.71 56.45 46.85 73.71 39.28 48.41 47.47 40.86
ViLT 35.32 48.24 47.85 62.70 56.44 58.06 60.22 73.66 34.62 42.41 44.43 59.51
CLIP 53.22 65.83 63.65 70.93 67.17 67.43 70.86 79.06 44.99 59.86 55.29 40.98
TMPT 55.41 61.61 67.67 76.60 63.19 67.25 62.92 81.19 43.56 59.24 55.68 46.82
TMPT+CoT 66.61 68.75 71.79 74.40 69.96 68.43 63.00 82.71 45.04 60.52 68.95 59.87
Qwen-VL 43.31 45.13 50.51 43.06 45.49 49.79 46.04 27.73 36.50 40.78 42.14 39.34
MiMo 66.21 65.89 61.54 45.32 42.19 50.69 42.36 40.26 39.30 58.95 60.07 52.39
LLaVa 49.48 45.80 60.30 20.48 28.48 33.83 38.94 31.57 36.53 47.47 50.45 47.07
Qwen2.5-VL 64.90 69.29 63.33 51.27 52.30 60.40 58.21 48.95 41.04 51.34 61.18 51.71
GPT4-Vision 70.46 72.82 61.63 44.59 47.07 57.47 57.90 37.61 44.83 56.40 66.72 56.90
[…]
Table 1: In-target multimodal stance detection results (%). Best and second-best scores are in bold and underline, respectively. indicates our ReMod model significantly outperforms baselines ($p<0.05$). The dashed line separates fine-tuned from non-fine-tuned methods. w/ denotes the backbone, and $\uparrow$ indicates improvement over the original MLLM.
Stance Inference.
During the inference phase, ReMoD first leverages the SAM to extract visual information $\mathcal{V}$ from the input image, capturing fine-grained visual manifestations of semantic entities. This visual information is then paired with the text input $u$ to form a multimodal context. It then employs its modality-synergistic retrieval mechanism to fetch the relevant prior experiences from both the modality and semantic experience pools. Finally, ReMoD integrates these experiences with multimodal context and the target to derive the final stance prediction, ensuring decisions are based on immediate data and guided by its refined long-term understanding.
Experiments
Experiment Settings
Dataset and Metrics.
To rigorously evaluate our ReMoD, we benchmark its performance on the MMSD dataset (1-liang2024multi), which comprises five distinct domains: Multi-modal COVID-CQ (MCCQ), Multi-modal Russo-Ukrainian Conflict (MRUC), Multi-modal Taiwan Question (MTWQ), Multi-modal Twitter Stance Election 2020 (MTSE), and Multi-modal Will-They-Won’t-They (MWTWT). Following the MMSD dataset, we adopt the Macro-F1 score as the primary metric for evaluation.
Implementation Details.
Our ReMoD framework is implemented upon the Qwen-2.5-VL-32B model as its backbone. For the experience pool retrieval, we set the coefficient $\alpha$ to 0.7, a relevance threshold $\tau$ to 0.8, and retrieve the top kk=3 experiences to balance context with noise. All input images are resized to a minimum of 28 pixels on their shorter side to meet requirements, while entity-related knowledge is sourced from Wikipedia using the Python Wikipedia library. More details are shown in the appendix.
Baseline Methods.
We evaluate our method against three categories of baselines: Text-only baselines: 1) BERT (33-devlin2019bert), using uncased BERT-base model; 2) RoBERTa (34-liu2019roberta), using RoBERTa-base model; 3) KEBERT (35-kawintiranon2022polibertweet); 4) LLaMA2 (36-touvron2023llama), using LLaMA2-70B-chat; 5) GPT4. Vision-only baselines: 1) ResNet (37-he2016deep), using ResNet-50 v1.5; 2) ViT (38-dosovitskiy2020image), using ViT-base-patch16-224 model; 3) Swin Transformer (SwinT) (39-liu2021swin), using SwinV2-base-patch4-window12-192-22k model. Multimodal baselines: 1) ViLT (40-kim2021vilt), using vilt-b32-mlm model; 2) CLIP (41-radford2021learning), using CLIP-ViT-base-patch32 model; 3) BERT+ViT, a model that uses BERT and ViT as the text and image encoders; 4) Qwen-VL (42-bai2023qwen), using the Qwen-VL-Chat7B. 5) Qwen2.5-VL (43-bai2025qwen2), using the Qwen2.5-VL-32B-Instruct; 6) GPT4-Vision ; 7) MiMo (44-xiaomi2025mimo), using MiMo-7B-RL, 8) LLaVa (45-li2024llava), using LLaVa-v1.6-Mistral-7B; 9) TMPT and TMPT+CoT(1-liang2024multi).
MODALITY METHOD MTSE MWTWT MRUC MTWQ
DT JB CA CE AC AH RUS UKR MOC TOC
Text-only BERT 32.52 29.97 63.55 61.30 59.18 52.89 22.01 15.45 28.04 9.57
RoBERTa 26.60 32.21 59.22 59.22 64.86 57.46 27.10 19.98 30.62 15.84
KEBERT 26.17 31.81 59.70 62.56 63.92 55.53 24.68 28.18 29.17 19.80
LLaMA2 53.57 53.92 32.47 38.37 48.08 46.13 31.86 36.34 51.46 44.10
GPT4 70.78 68.83 57.19 60.56 65.63 69.01 40.22 49.18 62.10 52.12
Vision-only ResNet 25.52 29.70 23.01 24.11 25.21 25.27 23.88 25.57 27.59 24.88
ViT 28.63 29.70 24.59 28.18 34.06 33.40 27.26 28.51 29.37 23.69
SwinT 28.54 30.85 28.53 28.50 35.87 34.33 25.44 24.54 27.90 19.69
Multimodal BERT+ViT 26.70 31.57 59.21 59.30 65.04 59.28 23.33 15.21 24.76 11.70
ViLT 28.08 29.74 38.33 46.00 55.01 48.55 21.56 23.96 23.54 19.18
CLIP 28.21 28.99 61.08 55.67 63.80 60.06 25.62 27.40 27.21 15.69
TMPT 31.69 32.65 66.36 66.39 66.32 61.56 23.87 24.71 32.18 26.48
TMPT+CoT 54.30 58.46 67.28 63.73 64.87 54.26 48.99 51.75 45.32 43.70
Qwen-VL 47.62 46.14 38.57 43.36 47.82 41.01 36.95 41.39 44.32 44.08
MiMo 66.58 66.01 38.11 40.10 44.20 39.05 44.85 50.24 64.65 46.25
LLaVa 50.26 46.74 26.53 30.77 34.96 36.57 40.67 38.69 48.97 39.80
Qwen2.5-VL 68.11 67.02 47.59 49.78 62.52 50.95 48.22 45.82 67.08 47.03
GPT4-Vision 72.68 71.28 42.23 45.92 54.59 53.19 42.09 47.00 65.00 52.36
[…]
Table 2: Zero-shot multimodal stance detection results (%). Best and second-best scores are in bold and underline, respectively. indicates our ReMod model significantly outperforms baselines ($p<0.05$). The dashed line separates fine-tuned from non-fine-tuned methods. w/ denotes the backbone, and $\uparrow$ indicates improvement over the original MLLM.
Main Results
In-target Multimodal Stance Detection.
As presented in Table 1, our ReMoD achieves competitive performance across most baselines, even powerful LLMs such as GPT4 and GPT4-Vision in multiple scenarios. While TMPT+CoT shows stronger results on some targets, it relies on training with GPT4-generated CoT examples. Instead, ReMoD highlights its training-free superiority by leveraging accumulated experience for adaptive modality contribution rethinking.
A deeper analysis of baseline performance reveals the intricate challenges in multimodal fusion. The consistent outperformance of text-only models over vision-only counterparts confirms that text serves as the primary modality for stance determination. Yet, the superior performance of advanced multimodal models like GPT4-Vision compared to their text-only variants validates the supplementary value of visual cues, underscoring the necessity of cross-modality integration. Notably, the inconsistent performance gains across fusion approaches, combined with the struggles of simpler frameworks like CLIP and TMPT, underscore a critical deficiency, as these models fail to handle the varying contribution of each modality on a per-sample basis, allowing noisy or irrelevant visual data to degrade overall performance. In contrast, ReMoD leverages its dual-reasoning paradigm to dynamically rethink modality contributions, thereby enhancing its final predictive accuracy.
Zero-shot Multimodal Stance Detection.
As shown in Table 2, the results expose the inherent fragility of models reliant on in-domain training. While smaller architectures like BERT and RoBERTa suffer a precipitous performance collapse, even powerful LLMs such as GPT4 and GPT4-Vision demonstrate significant performance volatility when confronted with unseen targets. This underscores a fundamental limitation in current paradigms: true zero-shot generalization remains a formidable and largely unsolved challenge. In this demanding scenario, ReMoD establishes a new standard for robustness and adaptability. It achieves state-of-the-art results on a majority of the evaluated datasets. This stability, particularly when contrasted with the erratic outputs of leading LLM-based baselines, provides compelling evidence of a superior generalization capability.
METHOD MTSE MCCQ MWTWT MRUC MTWQ
[…]
ReMoD 70.88 72.90 63.24 71.41 69.76
-w/o MEP 69.04 68.38 60.37 70.33 68.45
-w/o SEP 70.65 69.56 57.18 67.05 69.11
-w/o CoT 70.78 67.92 59.13 65.45 69.17
[…]
ReMoD 71.19 - 65.42 70.74 71.39
-w/o MEP 66.95 - 55.26 65.71 68.55
-w/o SEP 64.47 - 56.18 68.74 70.41
-w/o CoT 70.85 - 56.29 69.11 71.22
Table 3: Experimental results (%) of the ablation study on in-target and zero-shot multimodal stance detection.
Ablation Study
To evaluate the impact of each component within our ReMoD, we conduct an ablation study, with results reported in Table Zero-shot Multimodal Stance Detection.. The findings unequivocally demonstrate the synergistic importance of all modules, as the full ReMoD model consistently outperforms any ablated version across both in-target and zero-shot settings. The most critical components are the experience pools. Without using modality experience (w/o MEP) results in the most significant performance degradation, particularly in the zero-shot scenario, highlighting the indispensable role of assessing cross-sample modality contribution for robust generalization. Similarly, the exclusion of using semantic experience (w/o SEP) causes a severe decline, confirming the necessity of leveraging semantic context for accurate stance inference. Finally, while removing the CoT (w/o CoT) module also impairs performance, the impact is less pronounced, suggesting that CoT functions as a higher-level reasoning synthesizer, which organizes and refines the foundational evidence provided by the MEP and SEP to formulate a coherent final stance.
Figure 3: Results on different relevance thresholds.
Effect of Different Backbones
We assess the performance of ReMoD with different MLLM backbones in both in-target and zero-shot settings, as detailed in Table 1 and Table 2. Smaller backbones generally yield weaker performance, likely due to their limited visual perception capabilities. However, integrating any MLLM as ReMoD’s backbone still brings significant improvements marked by $\uparrow$ compared to standalone operation, with Qwen-VL showing particularly notable gains. This consistent enhancement across diverse architectures highlights that our framework effectively handles varying modality contributions by dynamically weighting them based on their actual expressive power for target stances. It thus compensates for limitations in smaller backbones, ensuring robust performance regardless of the backbone’s inherent biases toward text or vision.
Different Relevance Thresholds in Retrieval
We analyzed the effect of varying relevance thresholds (0.1, 0.3, 0.5, 0,7, 0.8, and 0.9) on retrieval performance for both in-target and zero-shot settings, as shown in Figure 3. The results indicate that in both settings, performance tends to be weaker at lower thresholds and gradually degrades once the threshold exceeds 0.8. This trend highlights a critical trade-off between the quality and quantity of retrieved experiences. A low threshold risks retrieving irrelevant or poor-quality experiences, which can introduce noise and impair the reasoning process. Conversely, an excessively high threshold becomes too restrictive, retrieving an insufficient number of experiences and thus limiting the model’s ability to leverage its accumulated knowledge for robust inference.
Figure 4: Case study between our ReMoD and directly CoT method based on the same LLM backbone.
Case Study
To further underscore the strengths of ReMoD’s dual reasoning, we analyze two examples in Figure 4 by comparing the results of direct CoT with ReMoD. In Example (a), direct CoT literally interprets ‘stan Putin’ in the text as support but ignores the negative context of ‘cringe’, which renders ‘stan’ sarcastic. ReMoD’s modality experience emphasizes weighting textual evidence via contextual judgment, enabling recognition of implicit sarcasm through the contradiction between cringe and stan to avoid being misled by isolated explicit terms. Its semantic experience clarifies that such sarcasm criticizes Russia-related phenomena, thus linking the text to the target stance of opposing Russia.
For Example (b), direct CoT infers support for the drug from surface information like India lifting the export ban in text and leaders’ collaboration in images. In contrast, ReMoD’s modality experience stresses prioritizing explicit stance information across modalities, equal weighting, and default neutrality, while it deepens understanding via common sense and geopolitical context, revealing that the text describes objective policy adjustments amid international games rather than the author’s subjective endorsement of the drug’s efficacy. By curbing surface information misinterpretation and calibrating reasoning with background knowledge, ReMoD accurately captures the neutral stance.
Conclusion
This paper proposes ReMoD, a framework for multimodal stance detection that employs a dual-reasoning paradigm inspired by human cognition. By integrating experience-driven intuitive reasoning with deliberate reflective reasoning, it dynamically captures the variable contributions of different modalities, overcoming the limitations of fixed fusion strategies. The dual-pool mechanism enables the accumulation and refinement of modality and semantic insights, enhancing robustness and context awareness. Experiments on MMSD show ReMoD outperforms most baselines in both in-target and zero-shot settings, with strong generalization.